Grade 12th pass9 grade science

The intial velocity of a body moving along a straight line is 7 ms−1 It has a uniform acceleration of 4ms−1. The distance covered by the body in the 5th second of its motion is :

Harshit Singh
4 Years agoGrade 12th pass

1 Answer

Yash Patari
4 Years ago
DEAR STUDENT

The distance covered by the body in nth second

Sn = u+ (a/2) (2n-1)

= 7+ (4/2) (2 x 5-1)

= 7+ 2(9)

= 7+ 18 = 25 m

THANKS